# Bayesian Multiobject Tracking With Neural-Enhanced Motion and Measurement Models

### v2

@@ V-B Feature Extraction and Computation of Learnable Factors
MSMDFusion[66] LiDAR+Camera 73.9 54.1 79.2 83.4 74.6 84.1 75.2 66.9 51.4 0.763 1088 743
FocalFormer3D-F [62] LiDAR+Camera 74.0 57.4 76.7 84.9 75.4 80.7 75.4 67.1 54.9 0.759 824 773
@@
By using kinematic features and shape features for each PO and measurement pair, $(i,j)$, we can compute the affinity factors $\mathpzc{f}^{i,j}_{\mathrm{af},k}$ for the proposed neural-enhanced measurement model. Specifically, for each $(i,j)$, the affinity factor $\mathpzc{f}^{i,j}_{\mathrm{af},k}$ is obtained as
$\mathpzc{f}^{i,j}_{\mathrm{af},k}=\mathrm{exp}\big{(}{\mathrm{NN}}\big{(}\underline{\bm{\mathbf{f}}}^{i}_{\mathrm{ki},k},\underline{\bm{\mathbf{f}}}^{i}_{\mathrm{sa},k},\overline{\bm{\mathbf{f}}}^{j}_{\mathrm{ki},k},\overline{\bm{\mathbf{f}}}^{j}_{\mathrm{sa},k}\big{)}\big{)}$ (40)
where $\mathrm{NN}(\cdot)$ is a neural network. The exponential mapping guarantees positivity and naturally controls the factor’s dynamic range compared to the corresponding model-based likelihood function $f(\bm{\mathbf{z}}^{j}_{k}|\underline{\bm{\mathbf{x}}}_{k}^{i})$. An optional rectified linear unit (ReLU) could be applied to $\mathpzc{f}^{i,j}_{\mathrm{af},k}$. This unit enforces that $\mathpzc{f}^{i,j}_{\mathrm{af},k}\geqslant 1$, i.e., the affinity factor can only increase the likelihood that measurement $j$ is associated with PO $i$ but not decrease it. It has been observed that applying the ReLU to $\mathpzc{f}^{i,j}_{\mathrm{af},k}$ can significantly reduce the training variance but potentially lead to a slightly reduced overall MOT performance.
The false positive rejection factor $f^{j}_{\mathrm{fpr},k}\in(0,1)$ for each measurement $j$, is computed by
$f^{j}_{\mathrm{fpr},k}=$ $\sigma\Big{[}{\mathrm{NN}}\big{(}\overline{\bm{\mathbf{f}}}^{j}_{\mathrm{ki},k},\overline{\bm{\mathbf{f}}}^{j}_{\mathrm{sa},k}\big{)}\Big{]}$ (41)
where $\sigma[\cdot]$ is the sigmoid function that restricts the output to $(0,1)$. Details on the loss function and training procedure are provided in the supplementary material [58].
VI Numerical Analysis – Scenario and Implementation Details
In this section, we will introduce the autonomous driving scenario used for performance evaluation. In particular, we employ LiDAR and camera data from the nuScenes autonomous driving dataset [18]. This dataset consists of 1000 autonomous driving scenes and seven object classes. Tracking is performed in 2-D. The kinematic state of objects is defined as $\bm{\mathbf{x}}^{i}_{k}=\big{[}(\bm{\mathbf{p}}^{i}_{k})^{\mathrm{T}}\hskip 0.85358pt(\bm{\mathbf{v}}^{i}_{k})^{\mathrm{T}}\big{]}^{\mathrm{T}}$, where $\bm{\mathbf{p}}^{i}_{k}\in\mathbb{R}^{2}$ and $\bm{\mathbf{v}}^{i}_{k}\in\mathbb{R}^{2}$ are the 2-D position and velocity, respectively. At each time step $k$, measurements are provided by a detector that consists of a neural network [62, 63]. In particular, for each detected object, measurements of the 2-D position, the 2-D velocity, a reliability score, as well as the size and orientation of the bounding box are provided. The 2-D position and 2-D velocity measurement, denoted as $\bm{\mathbf{z}}^{j}_{k}\in\mathbb{R}^{4}$ for each $j$, is directly used in the statistical model for MOT. In particular, the likelihood function $f(\bm{\mathbf{z}}^{j}_{k}|\bm{\mathbf{x}}_{k}^{i})$ used in (18) and (13) is based on the linear model
$\bm{\mathbf{z}}^{j}_{k}=\bm{\mathbf{x}}_{k}^{i}+\bm{\mathbf{r}}_{k}^{j}$ (42)
[…]
Measurements are pre-processed using the non-maximum suppression (NMS) technique [68]. Here, the threshold is set to 0.1, and no score filtering is performed. As discussed in Section IV-C, the prediction step is implemented using sigma points and results in a Gaussian representation of each PO state. The update step, on the other hand, is implemented using $10^{4}$ particles to represent each PO state. The Gaussian PDF resulting from the prediction step is used as a proposal PDF for the particle-based update step. From the particles resulting from the update step, a Gaussian representation is extracted by computing the sample mean and sample covariance. This Gaussian representation is then used as an input for the prediction step at the next time step. Using a particle-based implementation despite the linear sensor measurement model has been demonstrated to lead to superior data association results and thus an improved overall tracking performance [69].
Together with the probability of existence, the reliability score is used to compute the final AMOTA for each PO. The detection region of the sensor is given by $[p^{x}_{k}-54,p^{x}_{k}+54]\times[p^{y}_{k}-54,p^{y}_{k}+54]$, where $p^{x}_{k},p^{y}_{k}$ represent the time-dependent 2D positions of the ego vehicle. The prior PDF of false positives $f_{\mathrm{fp}}(\cdot)$ and newly detected objects $f_{\mathrm{n}}(\cdot)$ are uniformly distributed over the tracking region. The threshold for object declaration is $T_{\mathrm{dec}}=0.5$ and the threshold for pruning is $T_{\mathrm{pru}}=10^{-4}$ (cf. Section III-C). The survival probability is set to $p_{s}=0.999$. The parameters $p_{d}$, $\sigma^{2}_{p}$, $\sigma^{2}_{v}$, $Q$, $C_{\mathrm{n}}(\cdot)$, $C_{\mathrm{fp}}(\cdot)$ are estimated from the training data. are estimated from the training data.
The input data for the pretraining of the motion model is generated from ground truth information. In particular, to obtain the input data, random noise is added, deterministic biases are applied, and certain states are randomly removed from tracks. For the training of the measurement model, we add additional Gaussian noise to the measurements. This is motivated by the fact that the measurements generated by detectors in the training dataset are typically of higher quality than those in the validation dataset. The joint training of the proposed method is performed based on the Adam optimizer. The batch size and learning rate are set to 1 and $10^{-4}$. The tuning parameter $w_{\mathrm{\mathrm{fpr}}}$ in [58, Eq. (6)] is 0.1. The pre-training for the transition function is based on the Adam optimizer. The batch size and learning rate are set to 32 and $10^{-3}$. The following strategy has been identified as a means to make the training more stable. First, the motion model is pretrained individually, and then its training is refined by joint training with the measurement model. For the pretraining of the motion model, we can directly use ground-truth tracks and avoid the aforementioned assignment by means of the Hungarian algorithm. Details on feature extraction in the considered autonomous driving scenario are provided in the supplementary material [58].
@@ VI Numerical Analysis – Scenario and Implementation Details
TABLE II: Performance Using Different Detectors on Test Set
@@
VII Numerical Analysis – Real-Data Results
In what follows, we will present results in the considered autonomous driving scenario. In particular, we will present (i) an overall performance analysis for the proposed method, (ii) an ablation study for the proposed prediction step, and (iii) an ablation study for the update step.
VII-A Overall Performance Evaluation
The performance of the proposed NEMOT method is presented in Tables I and II. The rank of the different methods is determined based on the AMOTA. The results demonstrate that even that if combined with the “FocalFormer” detector [62], the proposed NEMOT method achieves the highest AMOTA and competitive AMOTP scores, compared to all other reference methods. A comparison across different methods based on Table I shows that, if both LiDAR and camera measurements are considered, our method achieves the best performance for trailers, truck, bicycle and pedestrian, and ranks second place for cars and motorcycles. Bicycles are notoriously challenging to distinguish from pedestrians in LiDAR point clouds. This insight explains the noticeable performance gap in AMOTA performance between methods that only rely on LiDAR data and methods that make use of LiDAR and camera data.
Also when combined with the “CenterPoint” detector [63], NEMOT outperforms all reference techniques. Additionally, as shown in Table II, NEMOT can achieve a lower IDS and Frag metric compared to methods reference methods including “ShaSTA” [40], “3DMOTFormer” [42] or “SimpleTrack” [35]. This is because the proposed method uses of a statistical model to determine the initialization and termination of tracks, which is more robust compared to the heuristic track management performed by these reference methods.
VII-B Ablation Study
TABLE III: Performance Contributed by Prediction and Update Enhancement on Validation Set
[…]
$\text{Prediction}(\times),\text{Update}(\times)$ 75.5 52.0
$\text{Prediction}(\times),\text{Update}(\checkmark)$ 76.7 49.5
$\text{Prediction}(\checkmark),\text{Update}(\times)$ 76.6 50.0
$\text{Prediction}(\checkmark),\text{Update}(\checkmark)$ 77.2 49.7
TABLE IV: Comparison of the Prediction MSE for Different Prediction Strategies
@@ VII-B Ablation Study
First, we evaluate the contributions of historical tracks and object interactions to prediction accuracy. The evaluation is performed using noisy versions of ground truth tracks as input for one-step prediction. Noisy versions of ground truth tracks are obtained by matching the closest measurement to ground truth tracks using the Hungarian algorithm. The resulting sequence of measurements is used as a noisy version of the ground truth track. The results, computed for different object types, are quantified using the MSE of the position by comparing the prediction results with the ground truth track. Using the noisy versions of ground truth track makes the prediction task more challenging and make performance differences more pronounced. This study compares approaches with three different motion models: (i) a baseline method that relies on a Markovian constant velocity (CV) motion model, (ii) a simplified version of the proposed method with a non-Markovian, learnable motion model which relies solely on the history of the to-be-predicted object state but does not include states of neighboring object states, and (iii) the proposed method with the complete non-Markovian, learnable motion model which relies on the history of the to-be-predicted object state and neighboring object states. The results shown in Table IV indicate the importance of using the history of the to-be-predicted object state and the history of neighboring states into account.
@@
TABLE V: Comparison of the AMOTA for The Enhanced Prediction in Tracking with Different Strategies
Method AMOTA $\uparrow$ AMOTP $\downarrow$
No SPs 76.1 51.6
Object SP Generation 76.6 50.2
Joint SP Generation 76.6 50.0
The results, presented in Table V, demonstrate that propagating uncertainty quantification by means of SP can significantly improve prediction performance. However, joint SP generation and separate SP generation achieve quite similar performance. Note that in our experiments, the separate strategy results in a runtime that is roughly reduced by 50 % compared to the joint strategy. The visualization of the improved prediction performance related to the proposed neural-enhanced motion model is provided in Fig. 5. An ablation study for the neural-enhanced measurement model is provided in the supplementary material [58].
We also conducted an ablation study for the neural-enhanced measurement model. Here, we compare the performance of the neural-enhanced update step that uses both affinity factors (denoted $\mathpzc{f}_{\mathrm{af}}$) and false alarm rejection factors (denoted $\mathpzc{f}_{\mathrm{fpr}}$), with neural-enhanced update steps that only use affinity factors and false alarm positive factors, respectively. It can be seen Table VI that the proposed approach yields the largest improvement in AMOTA performance only when both factors are considered. Both factors provide significant performance improvements. The positive impact of affinity factors and false positive rejection factors is further discussed and visualized in the supplementary material [58].
TABLE VI: Contributions of factors $\mathpzc{f}_{\mathrm{af}}$ and $\mathpzc{f}_{\mathrm{fpr}}$ in likelihood
@@
TABLE VII: Contributions of different types of features used for $\mathpzc{f}_{\mathrm{af}}$ computation
@@
Table VII, shows the contributions of each type of feature to factor $\mathpzc{f}_{\mathrm{af}}$. It can be seen that while only a subset of features already increases tracking performance compared to model-based MOT, the best result is achieved when the complete set of features is used. Another interesting insight is that even if only low-dimensional features are considered (i.e., position and velocity), this also leads to improved tracking performance. Notably, these low-dimensional features are already used within model-based MOT. This insight can be explained by the fact that the measurement model used by model-based MOT cannot accurately describe the underlying data-generating process, and only the proposed neural architecture can capture certain relationships of kinematic features with object states.

### v3

@@ VI-C Feature Extraction
MSMDFusion [11] LiDAR+Camera 74.0 57.6 76.7 84.9 75.4 80.7 75.4 67.1 54.9 0.763 1088 743
FocalFormer3D-F[19] LiDAR+Camera 73.9 54.1 79.2 83.4 74.6 84.1 75.2 66.9 51.4 0.759 824 773
@@
For object-oriented shape feature extraction, position, orientation, and size information determine the ROI $\underline{\mathcal{R}}_{1,k-1}^{i}$. The ROI is identified on the BEV feature map [68, 11, 69] in $\mathbb{R}^{180\times 180\times 128}$ corresponding to raw measurement from time k1k-1. Next, sample points from the ROI are determined using bilinear interpolation. In particular, we use edges or corners of the ROI. We apply two convolutional layers and a two-layer MLP to obtain the feature vector, i.e., $\mathrm{MLP}(\mathrm{Conv}(\underline{\mathcal{R}}_{1,k-1}^{i}))\hskip-0.85358pt\in\hskip-0.85358pt\mathbb{R}^{64}$. In addition, we perform an additional convolution over the entire BEV map and identify ROI $\underline{\mathcal{R}}_{2,k-1}^{i}$ on that resulting new BEV feature map in $\mathbb{R}^{180\times 180\times 64}$. By performing the same steps as above, we obtain a complementary feature vector $\mathrm{MLP}(\mathrm{Conv}(\underline{\mathcal{R}}_{2,k-1}^{i}))\hskip-0.85358pt\in\hskip-0.85358pt\mathbb{R}^{32}$. Finally, we concatenate these two feature vectors to obtain the final object-oriented shape feature, i.e.,
$\hskip-8.53581pt\underline{\bm{\mathbf{f}}}^{i}_{\mathrm{sa},k}=\big[\mathrm{MLP}(\mathrm{Conv}(\underline{\mathcal{R}}_{1,k-1}^{i}))^{\mathrm{T}}\hskip 2.84526pt\mathrm{MLP}(\mathrm{Conv}(\underline{\mathcal{R}}_{2,k-1}^{i}))^{\mathrm{T}}\big]^{\mathrm{T}}\hskip-0.85358pt\hskip-0.85358pt\hskip-0.85358pt\hskip-0.85358pt.$
The same feature-extraction procedure is performed to obtain measurement-oriented shape features $\overline{\bm{\mathbf{f}}}^{i}_{\mathrm{sa},k}$ using the BEV feature map corresponding to raw measurements from time $k$.
@@
Based on kinematic and shape features related to a PO $i$ and measurement $j$ pair, the affinity factor $\mathpzc{f}^{i,j}_{\mathrm{af},k}$ is obtained as
$\mathpzc{f}^{i,j}_{\mathrm{af},k}=$ $~\mathrm{exp}(\mathrm{MLP}(\Delta{\bm{\mathbf{f}}}^{i,j}))$
where we
introduced
$\Delta{\bm{\mathbf{f}}}^{i,j}=\Big[\big({\underline{\bm{\mathbf{p}}}}_{k}^{i}-{\overline{\bm{\mathbf{p}}}}_{k}^{j}\big)^{\hskip-0.85358pt\mathrm{T}}\hskip 2.84526pt{\underline{\bm{\mathbf{v}}}}^{i\hskip 0.85358pt\mathrm{T}}_{k}\hskip 2.84526pt{\underline{\bm{\mathbf{u}}}}^{i\hskip 0.85358pt\mathrm{T}}_{k}\hskip 2.84526pt{\underline{\bm{\mathbf{f}}}}^{i\hskip 0.85358pt\mathrm{T}}_{\mathrm{sa},k}\hskip 2.84526pt{\overline{\bm{\mathbf{v}}}}^{j\hskip 0.85358pt\mathrm{T}}_{k}\hskip 2.84526pt{\overline{\bm{\mathbf{u}}}}^{j\hskip 0.85358pt\mathrm{T}}_{k}\hskip 2.84526pt{\overline{\bm{\mathbf{f}}}}^{j\hskip 0.85358pt\mathrm{T}}_{\mathrm{sa},k}\Big]^{\hskip-0.85358pt\mathrm{T}}\hskip-0.85358pt\hskip-0.85358pt.$
Taking the difference of the position entries of the kinematic states removes the influence of the absolute values of position entries, which can vary strongly across scenes. The false positive rejection factor only relies on measurement-oriented features. In particular, for each measurement $j$, the false alarm rejection factor $\mathpzc{f}^{j}_{\mathrm{fpr},k}$ is obtained as
$\mathpzc{f}^{j}_{\mathrm{fpr},k}=$ $~\sigma\Big[{\mathrm{MLP}}({\overline{\bm{\mathbf{v}}}}^{j}_{k},{\overline{\bm{\mathbf{u}}}}^{j}_{k},\overline{\bm{\mathbf{f}}}^{j}_{\mathrm{sa},k})\Big].$
VII Numerical Analysis – Real-Data Results
In what follows, we will present results in the considered autonomous driving scenario. In particular, we will present (i) an overall performance analysis for the proposed method, (ii) an ablation study for the proposed prediction step, and (iii) an ablation study for the update step
@@ VII Numerical Analysis – Real-Data Results
TABLE II: Performance Evaluation Using Different Detectors Based on the Nuscenes Test Set
@@
VII-A Overall Performance Evaluation
The performance of the proposed NEMOT method is presented in Tables I and II. The ranking of the different methods is determined based on the AMOTA, which we consider the primary metric in our paper, consistent with the official NuScenes benchmark. The results demonstrate that even if combined with the “FocalFormer” detector [11], the proposed NEMOT method achieves the highest AMOTA and competitive AMOTP scores, compared to all other reference methods. A comparison across different methods based on Table I shows that, if both LiDAR and camera measurements are considered, our method achieves the best performance for trailers, truck, bicycle and pedestrian, and ranks second place for cars and motorcycles. Bicycles are notoriously challenging to distinguish from pedestrians in LiDAR point clouds. This insight explains the noticeable performance gap in AMOTA performance between methods that only rely on LiDAR data and methods that make use of LiDAR and camera data.
TABLE III: Performance Evaluation of Neural-Enhanced Prediction and Update Steps Based on Nuscenes Validation Set
[…]
$\text{Prediction}(\times),\text{Update}(\times)$ 75.5 52.0
$\text{Prediction}(\times),\text{Update}(\checkmark)$ 76.7 49.5
$\text{Prediction}(\checkmark),\text{Update}(\times)$ 76.6 50.0
$\text{Prediction}(\checkmark),\text{Update}(\checkmark)$ 77.2 49.7
Also when combined with the “CenterPoint” detector [69], NEMOT outperforms all reference techniques. Additionally, as shown in Table II, NEMOT can achieve a lower IDS and Frag metric compared to reference methods including “ShaSTA” [46], “3DMOTFormer” [65] or “SimpleTrack” [43]. This is because the proposed method uses a statistical model to determine the initialization and termination of tracks, which is more robust compared to the heuristic track management performed by these reference methods.
VII-B Ablation Study
@@
TABLE IV: Performance Comparison of Different Prediction Step Strategies
@@
TABLE V: Performance Comparison of Different Prediction Step Implementations
Method AMOTA $\uparrow$ AMOTP $\downarrow$ FPS $\uparrow$
No SPs 76.1 51.6 4.6
Object SP Generation 76.6 50.2 4.0
Joint SP Generation 76.6 50.0 2.0
(a) Ground Truth
(b) Estimation Result for $\mathpzc{f}_{\mathrm{af}}(\checkmark)$ and $\mathpzc{f}_{\mathrm{fpr}}(\times)$
(c) Estimation Result for $\mathpzc{f}_{\mathrm{af}}(\checkmark)$ and $\mathpzc{f}_{\mathrm{fpr}}(\checkmark)$
Figure 7: Tracking results for cars and pedestrians classes at one time step. The ground truth in (a) is compared with the proposed neural-enhanced update step. The neural enhanced update step is performed without false positive rejection (b) and with false positive rejection factors (c). Object declaration is performed by setting $T_{\mathrm{dec}}\hskip-0.85358pt=\hskip-0.85358pt0.7$. It can be seen that the result in (c) is closer to the result in (a) by having fewer false positives compared to (b).
(a) Likelihood Function – Three Cars and Clutter
(b) Likelihood Function – Two Close Pedestrians
Figure 8: Visualization of two examples that indicate how the affinity factors used by the neural-enhanced measurements update step can lead to improvements compared to a model-based measurement update step. Predicted particles and their corresponding weights provided by $l_{\mathrm{ne}}(\bm{\mathbf{z}}_{k}|\underline{\bm{\mathbf{y}}}^{i}_{k})$ and $l(\bm{\mathbf{z}}_{k}|\underline{\bm{\mathbf{y}}}^{i}_{k})$ are shown. A redder color indicates a higher particle weight. The black $\times$ denotes predicted position and the blue ++ denotes measurements. For this visualization, the false positive rejection factor is set to one. The neural-enhanced measurements update step assigns smaller weights to false positive measurements in (a) and more clearly distinguish between the two pedestrians in (b).
First, we evaluate the contributions of historical tracks and object interactions to prediction accuracy. The evaluation is performed using noisy versions of ground truth tracks as input for one-step prediction. Noisy versions of ground truth tracks are obtained by matching the closest measurement to ground truth tracks using the Hungarian algorithm. The resulting sequence of measurements is used as a noisy version of the ground truth track. The results, computed for different object types, are quantified using the MSE of the position by comparing the prediction results with the ground truth track. Using the noisy versions of the ground truth track makes the prediction task more challenging and makes performance differences more pronounced. This study compares approaches with three different motion models: (i) a baseline method that relies on a Markovian constant velocity (CV) motion model, (ii) a simplified version of the proposed method with learnable non-Markovian motion model which relies solely on the history of the to-be-predicted object state but does not include states of neighboring object states, and (iii) the proposed method with the complete learnable non-Markovian motion model which also make use of the history of the to-be-predicted object state and neighboring object states. The results shown in Table IV indicate the importance of using the history of the to-be-predicted object state and the history of neighboring states into account.
@@
The results, presented in Table V, demonstrate that propagating uncertainty quantification by means of SP can significantly improve prediction performance. However, joint SP generation and separate SP generation achieve quite similar performance. Table V also reports average Frames Per Second (FPS) to quantify the runtime of the three prediction strategies. FPS measurements have been obtained based on a laptop with a 13th Gen Intel(R) Core(TM) i9-13900HX CPU and an RTX 4060 GPU. Note that the separate strategy (ii) results in a runtime that is roughly reduced by 50 % compared to the joint strategy (iii). The visualization of the improved prediction performance related to the proposed neural-enhanced motion model is provided in Fig. 6.
We also conducted an ablation study for the neural-enhanced measurement model. Here, we compare the performance of the neural-enhanced update step that uses both affinity factors (denoted $\mathpzc{f}_{\mathrm{af}}$) and false alarm rejection factors (denoted $\mathpzc{f}_{\mathrm{fpr}}$), with neural-enhanced update steps that only use affinity factors and false alarm positive factors, respectively. It can be seen in Table VI that the proposed approach yields the largest improvement in AMOTA performance only when both factors are considered. Both factors provide significant performance improvements.
Table VII shows the contributions of each type of feature to factor $\mathpzc{f}_{\mathrm{af}}$. It can be seen that while only a subset of features already increases tracking performance compared to model-based MOT, the best result is achieved when the complete set of features is used. Another interesting insight is that even if only low-dimensional features are considered (i.e., position and velocity), this also leads to improved tracking performance. Notably, these low-dimensional features are already used within model-based MOT. This insight can be explained by the fact that the measurement model used by model-based MOT cannot accurately describe the underlying data-generating process, and only the proposed neural architecture can capture certain relationships of kinematic features with object states.
The positive impact of affinity factors is visualized in Fig. 8. Here, false positive rejection factors are set to one to make sure their effects are omitted. In particular, predicted particles weighted by the functions $l_{\mathrm{ne}}(\bm{\mathbf{z}}_{k}|\underline{\bm{\mathbf{y}}}^{i}_{k})$ and $l(\bm{\mathbf{z}}_{k}|\underline{\bm{\mathbf{y}}}^{i}_{k})$ used for model-based and neural-enhanced measurement update are shown. Fig. 8(a) depicts a scenario with three moving cars. Particles and their corresponding value of $l_{\mathrm{ne}}(\bm{\mathbf{z}}_{k}|\underline{\bm{\mathbf{y}}}^{i}_{k})$ and $l(\bm{\mathbf{z}}_{k}|\underline{\bm{\mathbf{y}}}^{i}_{k})$ for the three moving cars are shown. It can be observed that, contrary to the model-based measurement update, the neural-enhanced measurement update assigns lower particle weights to particles near false-positive measurements. Fig. 8(b) illustrates another potential advantage of the neural-enhanced update step in a scenario with two pedestrians. In particular, it can be seen that the neural-enhanced update step makes it possible to more clearly distinguish between the two pedestrians by assigning smaller weights to particles located in the area between the two pedestrians. Fig. 7, shows the positive effect of false positive rejection factors. In particular, it can be seen that the neural-enhanced update step with false positive rejection factors can clearly reduce the number of false positive object state estimates.
TABLE VI: Performance Evaluation – Contributions of Likelihood Factors $\mathpzc{f}_{\mathrm{af}}$ and $\mathpzc{f}_{\mathrm{fpr}}$
@@
$\mathpzc{f}_{\mathrm{af}}(\checkmark),\mathpzc{f}_{\mathrm{fpr}}(\times)$ 76.4 51.0
@@
TABLE VII: Performance Evaluation – Contributions of Different Types of Features Used for $\mathpzc{f}_{\mathrm{af}}$ Computation